02 Question (5 points) 1st attempt Part 1 (2 points) Feedback O See Hint A monopolist has a cost function given by c(y) = y2Part 3 (1 point) Feedback O See Hint * 100
Answer #1

pant 1 мл. I monopolist has a cost function ely)=43 and faes an inuuse de mard cave given by : pcy )= 120.00- y Now total rev20 120.00-2y = 24 -) -44 = - 120.00 : : ya .30.00 puce will ber Ply) = 120.00-y 2 120.00 – 30.00 . 90:00 output Huce, profitHunce What is you choose? Price ceiling = $ leve should 80.007 What quantity would be produced at this puce ecilings = 40.00